OpenStudy (anonymous):

-11,6 and the slope of 1

OpenStudy (aravindg):

can you please be more clear in your question ?

OpenStudy (anonymous):

A line that includes the point ( -11 , -6 ) has a slope of 1 . What is its equation in slope-intercept form?

OpenStudy (anonymous):

@AravindG

OpenStudy (aravindg):

can you take time to format it ??

OpenStudy (aravindg):

it looks like a direct copy paste from a test

OpenStudy (anonymous):

no test its a IXL.COM

OpenStudy (anonymous):

can you help ???????????????????????????

OpenStudy (anonymous):

@kailyn_lashon do u know about point slope form?

OpenStudy (anonymous):

a little but its just this question

OpenStudy (aravindg):

slope point form :: http://openstudy.com/updates/50f8cc3fe4b027eb5d9a556a

OpenStudy (anonymous):

A line that includes the point (-11,-6) has as slope of 1. What is its equation in slope-intercept form?

OpenStudy (anonymous):

@kailyn_lashon according to the point slope form, (y-y1) = m (x-x1) where m is the slope of the line and (x1 , y1) is any point on the line.

OpenStudy (anonymous):

using point form u will get the slope intercept form which is as y = mx + c where c is intercept